OpenStudy (0697169):

PLEASE HELP WILL GIVE MEDAL Simplify negative 6 and 1 over 4 – negative 2 and 1 over 8.

OpenStudy (0697169):

choices 4 and 1 over 8 negative 8 and 3 over 8 negative 4 and 1 over 8 8 and 3 over 8

OpenStudy (0697169):

@RANE

OpenStudy (0697169):

@Bobo-i-bo help plss

OpenStudy (bobo-i-bo):

ummm, is it: \[-6 \times \frac 1 4 - (-2) \times \frac 1 8\] ?

OpenStudy (0697169):

but none of the choiices say that

OpenStudy (mjdennis):

@0697169 Your question is not clear; could you restate it in math terms. If you cannot use the Equation editor that OpenStudy provides, could you: use * for multiplication and + for division and use / for division and - for subtraction

OpenStudy (mrnood):

\[-6\frac{ 1 }{ 4 } - (-2\frac{ 1 }{ 8 })\]

OpenStudy (mrnood):

first of all make both terms into imprpoer fractions with 8 as the denomitor e.g. \[2\frac{ 3 }{ 4 } = \frac{ 11 }{ 4 } = \frac{ 22 }{ 8 }\]